OpenStudy (anonymous):

Assume that y varies inversely as x. If y = -6 when x = -2, find y when x = 5.

OpenStudy (anonymous):

A. y = 1.7 B. y = 2.4 C. y = 3

OpenStudy (anonymous):

y=k/x -6=k/-2 -2*-6=k 12=k y=12/5

OpenStudy (anonymous):

IS IT 1.7?

OpenStudy (anonymous):

12/5 does not = 1.7

OpenStudy (anonymous):

5 goes into 12 how many times?

OpenStudy (anonymous):

2.4?

OpenStudy (vishal_kothari):

yes..